How many moles of Fe are needed to combine with 24 moles of Cl2 by this reaction? 




1
Expert's answer
2021-02-22T06:14:38-0500

2Fe + 3Cl2 = 2FeCl3

n(Fe) = "\frac{2}{3}"n(Cl2) = "\frac{2}{3} \times"24 = 16 (mol)

Answer: n(Fe) = 16 mol.